# Confidence-weighted integration of human and machine judgments for superior decision-making

### v2

@@ 4 Methods
Noisy LSVRC-2012 ImageNet [13]
A subset of the 2012 Large Scale Visual Recognition Challenge (LSVRC) ImageNet training set[17] was utilized. Namely, the dataset comprised 1 200 test cases (i.e., images), divided equally into 16 classes (chair, oven, knife, bottle, keyboard, clock, boat, bicycle, airplane, truck, car, elephant, bear, dog, cat, and bird). Two levels of phase noise were independently applied to distort the image dataset, low ($\Omega=80$) and high ($\Omega=125$).
@@
Procedure for human participants
Noisy LSVRC-2012 ImageNet [13]
One hundred forty-five participants classified between 34 and 74 noisy images (i.e., test cases) into 16 categories. For each evaluation, participants also provided a discrete confidence level (low, medium, or high). The total number of human classifications corresponded to 7 247 for low noise images, and 7 239 for high noise images.
BrainBench [14]
@@
Procedure for machines
Noisy LSVRC-2012 ImageNet [13]
Five different machine classifiers pretrained for ImageNet were utilized: AlexNet [18], DenseNet161 [19], GoogleNet [20], ResNet152 [21], and VGG-19 [22]. One pass through the noisy images data (epoch) was performed during stochastic gradient training. For a given image, the classifiers produced probability confidence scores for each of the 16 classes. The class label was assigned to the class with the highest probability confidence score.
BrainBench [14]
@@
Human and machine judgments were combined by adapting a Bayesian framework for human-machine complementarity [13]. The problem setting for combining two team members, human and machine, is as follows: Let $N$ denote the number of test cases to be analyzed with $L$ possible choices. The ground truth labels of the original abstracts are $z\in\{0,\ldots,L\}^{N}$. For the human classifier, the predicted labels $y\in\{0,\ldots,L\}^{N}$ and their corresponding confidence ratings $r\in\{0,\ldots,R-1\}^{N}$ with “0: lowest possible confidence” and “$R-1$: highest possible confidence” are given. For the machine classifier, we used the probability confidence scores $\pi\in\mathbb{R}_{+}^{N\times L}$. In the case of LLMs, $\pi=\text{Softmax}(-q)$, where $q$ is the PPL, reflecting a measure of uncertainty [14].
The first step of this model is to generate correlated probability confidence scores for human and machine classifiers using a bivariate normal distribution,
@@
$\mu_{H,i,j}$ $=b_{H}+(a_{H}-b_{H})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)\,,$
$\mu_{M,i,j}$ $=b_{M}+(a_{M}-b_{M})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)\,,$
with $\mathcal{Z}_{i}=\{x\,|\,x=z_{i}\}$. Note that the scalar parameters $a_{H}$, $a_{M}$, $b_{H}$, $b_{M}$, $\sigma_{H}$, $\sigma_{M}$, and $\rho_{HM}$ in Eq. (1) are learned from data. The parameter $\rho_{HM}$ from the covariance matrix learns the correlation between the human and machine classifiers. In the case of the machine classifier, $\pi_{M}$ is compared to the empirical probability confidence scores, $\pi$. Then, for the human classifier, $\pi_{H}$ is a latent variable that is used to calculate classifications
@@
where $\tau$ denotes a temperature parameter, usually small, that helps convergence [13]. The predicted classifications, $y_{H}$, are compared to actual human predictions, $y$. The confidence ratings are casted via an ordered logistic model [24] that maps the continuous probability confidence scores, $\pi_{H}$, to an ordinal confidence rating, $r_{H}$. This means,
$r_{H}\sim\text{OrderedLogistic}(\pi_{H},c,\delta)\,,$
@@
We introduce a logistic regression approach that combines the judgments of any number of teammates. The logistic combination model follows the principles of the Bayesian combination model, but is formulated within an easier to implement and extend regression framework. In its most basic form, which we consider here, each teammate is captured by a single predictor in the regression model. The value of the predictor on a trial depends on the teammate’s choice and their confidence. In particular, the magnitude of the predictor is the teammate’s confidence on that trial (i.e., confidence-weighted integration) and the sign is determined by the teammate’s choice. In general, the fitted beta weight for a teammate will reflect their accuracy and calibration. In a multiclass setting, multinomial regression is utilized.
As in the Bayesian combination model, $y$, $r$, $\pi$ are given. Thus, the logistic function for the $i$-th test case is of the form:
@@
$\beta^{\top}x=\beta_{I}+\beta_{H}\cdot x_{H}+\beta_{M}\cdot x_{M}\,.$ (2)
The fitted weights $\beta_{I}$, $\beta_{H}$, and $\beta_{M}$ correspond to the intercept, and human and machine teammates, respectively. The term $x_{k}$ is the signed confidence for team member $k$. For human participants, $x_{H}$ is $r_{i}$ if $y_{i}$ is the selected class label and $-r_{i}$ otherwise. Thus,
$x_{H}=r_{i}(1-2\cdot\mathbf{1}_{\mathcal{A}}(y_{i}))\,,$
where $\mathbf{1}_{\mathcal{A}}(y_{i})$ is an indicator function with $\mathcal{A}=\{x\,|\,x=1\}$. Hence, $\mathbf{1}_{\mathcal{A}}(x)=1$ if $x\in\mathcal{A}$, and $\mathbf{1}_{\mathcal{A}}(x)=0$ otherwise. For machine team members, it reads
$x_{M}=\pi_{i}(1-2\cdot\mathbf{1}_{\mathcal{A}}(y_{i}))\,.$
In the binary case of LLMs, we consider the PPL difference as
$x_{M}=|\Delta q_{i}|(1-2\cdot\mathbf{1}_{\mathcal{A}}(y_{i}))\,.$
@@
That more complex model should work better if an agent is a good predictor of outcomes but its confidence is random. In that case, $\alpha$ will be high, and the magnitude of the confidence weighting will always be $f(x)=1$ in Eq. (3). Thus, the fitted weights for logistic regression would not penalize the agent as much in the combination. For a really well-calibrated agent, $\alpha$ should be close to 0. Thus, Eq. (3) becomes the identity function, $f(x)=x$.
@@
Additionally, this model can be easily expanded by including additional terms to Eq. (2). For example, a third fitted weight could be included for an interaction term $x_{H}x_{M}$. Likewise, polynomial regression could be used to include $x_{H}^{2}$ and $x_{M}^{2}$, and corresponding fitted weights.
@@
Given $M$ observations for $N$ test cases (with $M>N$), we performed a leave-one-out cross-validation (LOOCV) [25, 26] that provides the best bias-variance trade-off for small datasets. Consider the evaluation of the $i$-th test case. With this procedure, we removed all instances of test case $i$ leaving the remaining $N-1$ test cases with all their instances to train the classifier teams. For testing, we utilized all the instances of test case $i$. This was repeated for all $N$ test cases, yielding $M$ predictions. Note that for individual teammates the evaluation comprised only the testing phase. When the predicted labels of any team or individual teammate in this study were randomly shuffled, the LOOCV accuracy dropped approximately to chance level (i.e., $1/L$).
@@ Acknowledgments
1:Set priors: $\alpha=1$, $a_{M}\sim\mathcal{N}(0,10)$, $b_{M}\sim\mathcal{N}(0,10)$, $\sigma_{M}\sim\text{Uniform}(0,15)$, $a_{H}\sim\mathcal{N}(0,10)$, $b_{H}=0$, $\sigma_{H}=1$, $\rho\sim\text{Uniform}(-1,1)$, $\tau=0.05$, $c\sim\text{Uniform}(0,1)$ with $c_{i}<c_{i+1}\;\forall\,i=1,\ldots,R-2$, and $\delta\sim\text{Uniform}(0,100)$.
2:start training:
3:     $p\sim\text{Dirichlet}(\alpha)$
4:     $z\leftarrow\text{argmax}(p)$$\triangleright$ Compare to actual data
5:     $\mu_{M,i,j}\leftarrow b_{M}+(a_{M}-b_{M})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
6:     $\mu_{H,i,j}\leftarrow b_{H}+(a_{H}-b_{H})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
7:     $\pi_{M}\sim\mathcal{N}\left(\mu_{M},\,\sigma_{M}\right)$triangleright Compare to actual data
8:     $\pi_{H}\sim\mathcal{N}\left(\mu_{H}+\rho\,\sigma_{H}\left(\frac{\pi_{M}-\mu_{M}}{\sigma_{M}}\right),\,\sqrt{1-\rho^{2}}\,\sigma_{H}\right)$
9:     $y\sim\text{Categorical}\left(\text{Softmax}\left(\pi_{H}/\tau\right)\right)$triangleright Compare to actual data
10:     $r\sim\text{OrderedLogistic}(\pi_{H},c,\delta)$triangleright Compare to actual data
11:end training
12:start testing:
13:     $p\sim\text{Dirichlet}(\alpha)$
14:     $z\leftarrow\text{argmax}(p)$
15:     $\mu_{M,i,j}\leftarrow b_{M}+(a_{M}-b_{M})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
16:     $\mu_{H,i,j}\leftarrow b_{H}+(a_{H}-b_{H})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
17:     $\pi_{M}\sim\mathcal{N}\left(\mu_{M},\,\sigma_{M}\right)$triangleright Compare to actual data
18:     $\pi_{H}\sim\mathcal{N}\left(\mu_{H}+\rho\,\sigma_{H}\left(\frac{\pi_{M}-\mu_{M}}{\sigma_{M}}\right),\,\sqrt{1-\rho^{2}}\,\sigma_{H}\right)$
19:     $y\sim\text{Categorical}\left(\text{Softmax}\left(\pi_{H}/\tau\right)\right)$triangleright Compare to actual data
20:     $r\sim\text{OrderedLogistic}(\pi_{H},c,\delta)$triangleright Compare to actual data
21:end testing
Algorithm S1 illustrates the inference procedure of the Bayesian combination model, which comprises two stages: estimation of parameters (training) and prediction of label probabilities (testing). For the estimation of parameters, the given data is: true labels $z$, probability confidence scores $\pi_{M}$, and human classification $y$ and confidence ratings $r$. We assumed that all human participants shared the same set of parameters ($a_{H}$ , $b_{H}$ , $\sigma_{H}$, $c$, $\delta$, and $\tau$). In the neuroscience forecasting task, human confidence ratings on the slider bar were mapped to range between 1 and 100. A wide range is computationally expensive, thus, we aggregate it into three levels: “0: low confidence”, “1: moderate confidence”, and “2: high confidence”. Then, the aggregated confidence rating used for analysis, $r\in\{0,1,2\}^{N}$, reads
@@
The utilized cutpoints (i.e., 33 and 66) produced a good agreement between confidence and accuracy. Among the evaluations of human participants, “low” had $63.2\%$ average accuracy ($n=174$), “moderate” had $66.5\%$ ($n=185$), and “high” had $81.9\%$ ($n=144$).
To estimate the posterior over the underlying parameters, a No-U-Turn Sampler (NUTS) for Markov chain Monte Carlo (MCMC) [27] was used with $n_{w}=500$ warmup steps, $n_{c}=3$ chains, and $n_{s}=25$ samples. For the prediction of label probabilities the given data is the same as before but $z$, that is now latent. The prior distributions were set according to [13]. The correlation between team members was $\rho\sim\mathrm{Uniform}(-1,1)$. For the LLM classifier, $a\sim\mathcal{N}(0,10)$, $b\sim\mathcal{N}(0,10)$, and $\sigma\sim\mathrm{Uniform}(0,15)$. For human classifications, $a$ was the same but $b$ and $\sigma$ were adjusted to $b=0$ and $\sigma=1$ for the purpose of identifiability [13]. In addition, we used $\delta\sim\mathrm{Uniform}(0,100)$ for the scaling parameter and uniform priors on the cutpoints, $c\sim\mathrm{Uniform}(0,1)$, with the constraint that the cutpoints are ordered (i.e. $c_{i}<c_{i+1}\;\forall\,i=1,\ldots,R-2)$. Finally, we set $\tau=0.05$ for best convergence results [13].
@@
Figure S1: Performance of the Bayesian and confidence-weighted logistic combination models in a low noise object recognition task [13]. Accuracy results on low levels of image noise ($\Omega=80$) with the Bayesian and confidence-weighted combination models, where humans outperform machines ($t(4)=-4.77$, $P<0.01$). Each data point corresponds to the average across 7 247 image evaluations. Error bars represent standard error of the mean using a binomial model. The performance between the Bayesian and confidence-weighted combination models is indistinguishable for human-machine teams ($t_{Welch}(6)=0.49$, $P=0.32$). However, the confidence-weighted combination model presented improved results for machine-machine teams ($t_{Welch}(16)=3.75$, $P<0.001$).
@@
Figure S2: Removing confidence from the confidence-weighted logistic combination model does not negatively impact team performance in a noisy object recognition task [13]. Accuracy results on high levels of image noise ($\Omega=125$) with the confidence-weighted logistic combination model, where confidence scores were set to 1, i.e., $f(x)=1$ in Eq. (3). Each data point corresponds to the average across 7 239 image evaluations. Error bars represent standard error of the mean using a binomial model. Similarly as in the standard scenario (Figure 1), human-machine teams surpass machine-only teams ($t_{Welch}(24)=2.80$, $P<0.01$), although with a smaller margin.
@@
Figure S3: Optimal squashing in the confidence-weighted logistic combination model does not improve overall team performance in the neuroscience forecasting task [14]. Accuracy results on the neuroscience forecasting task with the confidence-weighted logistic combination model, where the parameter $\alpha$ in Eq. (3) is optimized. Llama2 chat 7B, 13B, and 70B models were considered. Each data point corresponds to the average across 503 test case evaluations. Error bars represent standard error of the mean using a binomial model. Similarly as in the standard scenario (Figure 4), human-LLM teams surpass LLM-only teams ($t_{Welch}(11)=7.20$, $P<0.0001$). The performance between standard and squashing scenarios is, however, indistinguishable ($t_{Welch}(19)=0.05$, $P=0.48$).
@@
Figure S4: Adding interaction terms to the confidence-weighted logistic combination model does not improve overall team performance in the neuroscience forecasting task [14]. Accuracy results of the confidence-weighted logistic combination model on the neuroscience forecasting task as a function of the degree of the polynomial expansion of the features. Llama2 chat 7B, 13B, and 70B models were considered. Each data point corresponds to the average across 503 test case evaluations. Only interactions were considered, as the results including pure terms is identical. Team performance is independent of interaction terms.

### v3

@@ 4 Methods
ImageNet 16H [13]
A subset of the 2012 Large Scale Visual Recognition Challenge (LSVRC) ImageNet training set [20] was utilized. Namely, the dataset comprised 1 200 test cases (i.e., images), divided equally into 16 classes (chair, oven, knife, bottle, keyboard, clock, boat, bicycle, airplane, truck, car, elephant, bear, dog, cat, and bird). Four levels of phase noise were independently applied to distort the image dataset ($\Omega=\{80,95,110,125\}$). We considered two noise levels: low ($\Omega=80$), where humans outperform machines (Figure S2), and high ($\Omega=125$), where most machines outperform humans (Figure 1).
One hundred forty-five participants classified between 34 and 74 noisy images (i.e., test cases) into the 16 aforementioned categories. For each evaluation, participants also provided a discrete confidence level (low, medium, or high). The total number of human classifications corresponded to 7 247 for low noise images, and 7 239 for high noise images.
Five different machine classifiers pretrained for ImageNet 16H were utilized: AlexNet [21], DenseNet161 [22], GoogleNet [23], ResNet152 [24], and VGG-19 [25]. One pass through the noisy images data (epoch) was performed during stochastic gradient training. For a given image, the classifiers produced probability scores for each of the 16 classes. The class label was assigned to the class with the highest probability.
@@
Human and machine judgments were combined by adapting a Bayesian framework for human–machine complementarity [13]. The problem setting for combining two team members, human and machine, is as follows: Let $N$ denote the number of test cases to be analyzed with $L$ possible choices. The ground truth labels of the original test cases are $z\in\{0,\ldots,L-1\}^{N}$. For the human classifier, the predicted labels $y\in\{0,\ldots,L-1\}^{N}$ and their corresponding confidence ratings $r\in\{0,\ldots,R-1\}^{N}$ with “0: lowest possible confidence” and “R1R-1: highest possible confidence” are given. For the machine classifier, we used the probability scores $\pi\in\mathbb{R}_{+}^{N\times L}$. In the case of LLMs, $\pi=\text{Softmax}(-q)$, where $q$ are the PPL scores, reflecting a measure of uncertainty [14].
The first step of this model is to generate correlated probability scores for human and machine classifiers using a bivariate normal distribution,
@@
$\mu_{i,j,H}$ $=b_{H}+(a_{H}-b_{H})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)\,$,
$\mu_{i,j,M}$ $=b_{M}+(a_{M}-b_{M})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)\,$,
with $\mathcal{Z}_{i}=\{x\,|\,x=z_{i}\}$. Note that the scalar parameters $a_{H}$, $a_{M}$, $b_{H}$, $b_{M}$, $\sigma_{H}$, $\sigma_{M}$, and $\rho_{HM}$ in Equation (1) are learned from data. The parameter $\rho_{HM}$ from the covariance matrix learns the correlation between the human and machine classifiers. In the case of the machine classifier, $\pi_{M}$ is compared to the empirical probability scores, $\pi$. Then, for the human classifier, $\pi_{H}$ is a latent variable that is used to calculate classifications
@@
where $\tau$ denotes a temperature parameter, usually small, that helps convergence [13]. The predicted classifications, $y_{H}$, are compared to actual human predictions, $y$. Ordered logit and probit models [27] yield practically indistinguishable results [28]. Departing from Steyver’s et al.’s [13], we used ordered logit because of software availability. We successfully reproduced their prior results with probit using ordered logit (see Figure S1). Ordered logit maps the continuous probability scores, $\pi_{H}$, to an ordinal confidence rating, $r_{H}$. This means,
$r_{H}\sim\text{OrderedLogistic}(\pi_{H},c,\delta)\,$, (2)
@@
We introduce a logistic regression approach that combines the judgments of any number of teammates. The logistic combination model follows the principles of the Bayesian combination model, but is formulated within an easier to implement and extend regression framework. In its most basic form, which we consider here, each teammate is captured by a single predictor in the regression model. The value of the predictor on a trial depends on the teammate’s choice and their confidence. In particular, the magnitude of the predictor is the teammate’s confidence on that trial (i.e., confidence-weighted integration) and the sign is determined by the teammate’s choice. In general, the fitted beta weight for a teammate will reflect their accuracy and calibration.
As in the Bayesian combination model, $y$, $r$, $\pi$ are given. The logistic function is of the form:
@@
$\beta^{\top}x=\beta_{I}+\beta_{H}x_{H}+\beta_{M}x_{M}\,.$ (3)
The fitted weights $\beta_{I}$, $\beta_{H}$, and $\beta_{M}$ correspond to the intercept, and human and machine teammates, respectively. The term $x_{i,j,k}$ is the signed confidence of the $i$-th test case for the class label $j$ selected by team member $k$. For human participants, $x_{i,j,H}$ is $r_{i}$ if $y_{i}=j$ is the selected class label and $r_{i}$ otherwise. Similarly, for machine team members, $x_{i,j,M}=\pi_{i,j}$ if $y_{i}=j$ is the selected class label and $x_{i,j,M}=-\pi_{i,j}$ otherwise. In the binary case of LLMs, we consider the absolute PPL difference as a measure of confidence. This means, $x_{i,0,M}$ is $|\Delta q_{i}|$ if $y_{i}=0$ is the selected class label and $|\Delta q_{i}|$ otherwise.
@@
That more complex model should work better if an agent is a good predictor of outcomes but its confidence is random. In that case, $\alpha$ will be high, and the magnitude of the confidence weighting will always be $f(x)=1$ in Equation (4). Thus, the fitted weights for logistic regression would not penalize the agent as much in the combination. For a really well-calibrated agent, $\alpha$ should be close to 0. Thus, Equation (4) becomes the identity function, $f(x)=x$.
@@
Additionally, this model can be easily expanded by including additional terms to Equation (3). For example, a third fitted weight could be included for an interaction term $x_{H}x_{M}$. Likewise, polynomial regression could be used to include $x_{H}^{2}$ and $x_{M}^{2}$, and corresponding fitted weights.
@@
Given a total of $M$ observations across $N$ different test cases (with $M>N$), we performed a leave-one-out cross-validation (LOOCV) that provides the best bias-variance trade-off for small datasets. We evaluated the consistency of performance estimates across different cross-validation procedures, demonstrating that our results reflect true model generalization rather than artifacts of a particular validation strategy (Figure S8).
Consider the evaluation of the $i$-th test case. With this procedure, we removed all instances of test case $i$ leaving the remaining N1N-1 test cases with all their instances to train the classifier teams. For testing, we utilized all the instances of test case $i$. This was repeated for all $N$ test cases, yielding $M$ predictions. Note that for individual teammates the evaluation comprised only the testing phase. When the predicted labels of any team or individual teammate in this study were randomly shuffled, the LOOCV accuracy dropped approximately to chance level (i.e., $1/L$).
Data and code availability
For ImageNet 16H [13] and BrainBench [14], the previously reported human participant data and machine confidence scores utilized in each study are available at https://osf.io/2ntrf and https://github.com/braingpt-lovelab/BrainBench, respectively. All computer code associated with this work including combination model implementations, team evaluations, and analyses are publicly available at https://github.com/braingpt-lovelab/haico and archived in Edmond [29].
@@ Acknowledgments
We implemented the Bayesian combination model [13] in Python, which was originally developed in JAGS. The model comprises two stages: parameter inference (training) and class label prediction (testing). Algorithm S1 illustrates the inference procedure for the human–machine case. During testing, the posterior samples of the parameters learned in training are used to evaluate unseen data. For each posterior sample, we compute the joint log-likelihood of the machine scores, human classifications, and human confidence ratings under each possible class. These log-likelihoods are aggregated across samples and normalized via a softmax transformation. The predicted class label corresponds to the class with the highest probability.
In Figure S1, we reproduced the experiment from Figure 3 (Top) in Steyver’s et al. [13] using our Python implementation. The overall team performance ($n=16$) obtained with our method is indistinguishable from the original JAGS implementation (Welch’s t(28.95)=0.24t(28.95)=-0.24, $P=0.81$). This also holds for the five human–machine (Welch’s t(5.33)=2.17t(5.33)=-2.17, $P=0.08$), and ten machine–machine (Welch’s $t(17.60)=0.17$, $P=0.87$) teams.
@@
1:Given data: true class labels $z$, probability scores $\pi_{M}$, human classification $y$, and confidence ratings $r$.
2:Set priors: $a_{M}\sim\mathcal{N}(0,10)$, $b_{M}\sim\mathcal{N}(0,10)$, $\sigma_{M}\sim\text{Uniform}(0,15)$, $a_{H}\sim\mathcal{N}(0,10)$, $b_{H}=0$, $\sigma_{H}=1$, $\rho\sim\text{Uniform}(-1,1)$, $\tau=0.05$, $c\sim\text{Uniform}(0,1)$ with $c_{i}<c_{i+1}\;\forall\,i=1,\ldots,R-2$, and $\delta\sim\text{Uniform}(0,100)$.
3:start parameter inference:
4:  $\mu_{i,j,M}\leftarrow b_{M}+(a_{M}-b_{M})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
5:  $\mu_{i,j,H}\leftarrow b_{H}+(a_{H}-b_{H})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
6:  $\pi_{M}\sim\mathcal{N}\left(\mu_{M},\,\sigma_{M}\right)$ $\triangleright$ Compare to actual data
7:  $\pi_{H}\sim\mathcal{N}\left(\mu_{H}+\rho\,\sigma_{H}\left(\frac{\pi_{M}-\mu_{M}}{\sigma_{M}}\right),\,\sqrt{1-\rho^{2}}\,\sigma_{H}\right)$
8:  $y\sim\text{Categorical}\left(\text{Softmax}\left(\pi_{H}/\tau\right)\right)$ $\triangleright$ Compare to actual data
9:  $r\sim\text{OrderedLogistic}(\pi_{H},c,\delta)$ $\triangleright$ Compare to actual data
10:end parameter inference
We assumed that all human participants shared the same set of parameters ($a_{H}$ , $b_{H}$ , $\sigma_{H}$, $c$, $\delta$, and $\tau$). In the neuroscience forecasting task, human confidence ratings on the slider bar were mapped to range between 1 and 100. A wide range is computationally expensive, thus, we aggregate it into three levels: “0: low confidence”, “1: moderate confidence”, and “2: high confidence”. Then, the aggregated confidence rating used for analysis, $r\in\{0,1,2\}^{N}$, reads
@@
The utilized cutpoints (i.e., 33 and 66) produced a good agreement between confidence and accuracy. Among the evaluations of human participants, “low” had $63.2\%$ average accuracy ($n=174$), “moderate” had $66.5\%$ ($n=185$), and “high” had $81.9\%$ ($n=144$). To infer the posterior over the underlying parameters, a No-U-Turn Sampler (NUTS) for Markov chain Monte Carlo (MCMC) [30] was used with $n_{w}=1000$ warmup steps, $n_{c}=8$ chains, and $n_{s}=50$ samples.
Figure S1: Performance of the Bayesian combination models in the low-noise object recognition task [13].
We reproduced Steyver’s et al.’s [13] results using ordered logit (see Equation 2), as opposed to probit. Accuracy results on low levels of image noise ($\Omega=80$) with the Bayesian combination model, reproducing the original setting. We used 1 200 predictions (corresponding to 1 200 unique images) and split them into four random partitions for the purpose of four-fold cross-validation. Two human participants were created by random sampling without replacement across the 1 200 unique images. In this case, machine classifiers are surpassed by humans (t(4)=3.88t(4)=-3.88, $P<0.01$). Teams comprising a human and a machine (green points) deliver superior results compared to machine-only and machine–machine teams (Welch’s $t(16.04)=9.18$, $P<0.0001$). Each data point corresponds to the average across 1 200 image evaluations. Error bars represent standard error of the mean using a binomial model.
Figure S2: Performance of the confidence-weighted logistic combination model in the low-noise object recognition task [13].
@@
Accuracy results on low levels of image noise ($\Omega=80$) with the confidence-weighted combination model, where humans outperform machines (t(4)=4.77t(4)=-4.77, $P<0.01$). Human–machine teams consistently outperform teams with one or more machines (Welch’s $t(37.46)=12.50$, $P<0.0001$). Each data point corresponds to the average across 7 247 image evaluations. Error bars represent standard error of the mean using a binomial model.
Figure S3: Conditions for effective collaboration between human experts and machines are satisfied for the noisy object recognition task [13].
(A) When human experts and machines were confident in their judgments, they were more likely to be correct. Confidence ratings were sorted into equal bins, and the mean accuracy for each bin was plotted. The positive slope of the black regression lines for humans and models (AlexNet, DenseNet161, GoogleNet, ResNet152, and VGG-19) indicates well-calibrated confidence [14, 17, 18, 19], meaning higher confidence correlates with higher accuracy.
(B) Item difficulty Spearman correlations among machines and human experts. For machines, we used the probability score of the predicted image class to determine relative image classification difficulty. The probability score was then signed with +1+1 if the prediction was correct and 1-1 if it was incorrect. Mean accuracy was used for human experts. Heatmap color scale ranges from 0.1 to 0.9.
(C) Besides AlexNet, models surpass human experts on this task. Error bars represent standard error of the mean using a binomial model.
Figure S4: Removing confidence from the confidence-weighted logistic combination model does not negatively impact team performance in the noisy object recognition task [13].
@@
Accuracy results on high levels of image noise ($\Omega=125$) with the confidence-weighted logistic combination model, where the probability scores were set to 1, i.e., $f(x)=1$ in Equation 4. Similarly as in the base scenario (Figure 1), human–machine teams surpass machine-only teams (Welch’s $t(24.08)=2.80$, $P<0.01$), although with a smaller margin. Each data point corresponds to the average across 7 239 image evaluations. Error bars represent standard error of the mean using a binomial model.
Figure S5: Confidence assessment in the noisy object recognition task [13].
Histogram of top-1 probability scores under high image noise ($\Omega=125$) for different machine classifiers. For each test case ($n=1200$), only the highest predicted probability was used, and values were grouped into 30 bins. The resulting distributions are heavily skewed toward 1. To test whether confidence magnitudes improve predictions, we used cross-entropy, a metric that reflects both accuracy and calibration. For each classifier, cross-entropy was computed independently for both signed probability scores (i.e., signed confidence) and sign-only features. The results are nearly identical (Welch’s $t(7.99)=0.24$, $P=0.82$), suggesting that confidence magnitudes provide no additional benefit in this setting.
Figure S6: Optimal squashing in the confidence-weighted logistic combination model does not improve overall team performance in the neuroscience forecasting task [14].
@@
Accuracy results on the neuroscience forecasting task with the confidence-weighted logistic combination model, where the parameter $\alpha$ in Equation 4 was optimized. Similarly as in the base scenario (Figure 4), human–LLM teams surpass LLM-only teams (Welch’s $t(11.89)=7.20$, $P<0.0001$). The performance between base and squashing scenarios is, however, indistinguishable (Welch’s $t(19.32)=0.05$, $P=0.96$). Each data point corresponds to the average across 503 test case evaluations. Error bars represent standard error of the mean using a binomial model.
Figure S7: Adding interaction terms to the confidence-weighted logistic combination model does not improve overall team performance in the neuroscience forecasting task [14].
@@
Accuracy results of the confidence-weighted logistic combination model on the neuroscience forecasting task as a function of the degree of the polynomial expansion of the features. Only interaction terms were considered in the polynomial expansion, as the results including pure terms were identical. Team performance is independent of interaction terms. Each data point corresponds to the average across 503 test case evaluations.
Figure S8: Performance consistency analysis across different cross-validation procedures.
Box plots show the distribution of overall accuracy for 3-fold, 5-fold, 10-fold, 50-fold, and leave-one-out (LOO) cross-validation (CV) procedures for the experiment presented in Figure 4. Overlaid data points represent human–machine and machine–machine team predictions. Paired t-tests (parametric) and Wilcoxon signed-rank tests (non-parametric) were used to compare each CV procedure with LOOCV to ensure robust assessment regardless of data distribution assumptions. Results show non-significant differences, with similar mean accuracies (ranging from 77.0% to 78.7%) and overlapping 95% bootstrap confidence intervals (widths between 1.8% and 3.2%). This demonstrates that LOOCV is an appropriate validation procedure to evaluate model performance.